1 person can paint a house in 5 hours another person can paint the same house in 3 hours and another person can paint the house in 2 hours how long will it take all three people working together to paint the house?
--------------------------------------------------------------------------------

Answer by mananth(16949)

1 person can paint a house in 5 hours another person can paint the same house in 3 hours and another person can paint the house in 2 hours how long will it take all three people working together to paint the house?

A------------ 5 hours------------- 1/5 of job in 1 hour
B-------------3 hours--------------1/3 of the job in 1 hour
C-------------2 hours--------------1/2 of the job in 1 hour
..
TOGETHER They do
1/5 + 1/3 + 1/2
= (6+10+15)/30
31/30 of thejob in 1hour=
So they will take 31/30 hours = 1.03 hours

m.ananth@hotmail.ca